# Infinity-Parser: Layout-Aware Reinforcement Learning for Scanned Document Parsing

### v1

@@ 2.2 RL with Layout-Aware Rewards
As illustrated in Figure 4, we employ a RL framework to directly optimize scanned document parsers, aiming to enhance both structural fidelity and semantic accuracy. To achieve this, we utilize GRPO [42], which enables learning from rule-based reward signals without relying on absolute values.
GRPO operates by generating a set of candidate Markdown outputs for each document and evaluating them using a multi-aspect reward, denoted as $R_{\text{Multi-Aspect}}$, which integrates multiple rule-based criteria into a unified supervisory signal. These raw rewards are then converted into relative advantage scores by comparing each candidate against others within the same group. This relative evaluation promotes training stability and encourages the selection of higher-quality outputs, eliminating the need for a learned value function or critic.
@@
1. Edit Distance Reward ($R_{\mathrm{dist}}$)
We define the edit distance reward based on the normalized Levenshtein distance $D(y,\hat{y})$ between the predicted output $\hat{y}$ and the reference output $y$:
$R_{\text{dist}}=1-\frac{D(y,\hat{y})}{\max(N,M)},$
where $N=|y|$ and $M=|\hat{y}|$ are the lengths of the reference and predicted sequences, respectively. The distance $D(y,\hat{y})$ measures the minimum number of single-character insertions, deletions, or substitutions required to convert $\hat{y}$ into $y$, thereby capturing both semantic and formatting discrepancies. This reward is bounded within $[0,1]$, with higher values indicating better alignment between prediction and reference. Meanwhile, the reference output $y$ is synthesized through two data generation pipelines proposed in this work. It is produced with rigorous rule-based filtering and consistency validation, and serves as a high-quality surrogate for ground-truth annotations in evaluating the quality of the model output.
2. Count Reward ($R_{\mathrm{count}}$)
@@
$R_{\text{count}}=1-\frac{|N_{Y}-N_{\hat{Y}}|}{N_{Y}},$
@@
3. Order Reward ($R_{\mathrm{order}}$)
@@
$R_{\text{order}}=1-\frac{D_{\text{order}}}{\max_{\text{inv}}},$
@@
The final multi-aspect reward is a weighted combination of these three components. Specifically, we begin by applying the Hungarian algorithm [20] to establish the optimal one-to-one matching between predicted and ground-truth segments, identifying both pairings and their relative order. Based on the matched segment count, we compute the count reward to reflect alignment in the number of segments. Using the relative sequence of matched pairs, we calculate the order reward to measure structural consistency. On top of these matchings, we compute the edit reward by averaging the edit similarities of each matched segment pair. Combining these terms yields the final reward:
$R_{\text{Multi-Aspect}}=R_{\text{dist}}+R_{\text{count}}+R_{\text{order}}$
@@ 3 Experiments
We fine-tune the Qwen2.5-VL-7B model using GRPO within a distributed training setup based on Verl [45, 66], utilizing 8 A100 GPUs (80GB). Throughout our experiments, we set the KL coefficient $\beta=1.0\times 10^{-2}$. And for each problem instance, we sample 8 responses, each with a maximum length of 8192 tokens and a temperature of 1.0. Both the rollout batch size and the global batch size are set to 128. The actor model is updated using the AdamW optimizer with parameters $(\beta_{1}=0.9,\beta_{2}=0.99)$ and a learning rate $1.0\times 10^{-6}$. The model is trained for 1.0 epoch for all experiments.
[…]
EN ZH EN ZH EN ZH EN ZH EN ZH EN ZH EN ZH
Based on Pipeline Tools
MinerU [52] 0.061 0.215 0.278 0.577 57.3 42.9 78.6 62.1 0.18 0.344 0.079 0.292 0.15 0.357
Marker  [37] 0.080 0.315 0.530 0.883 17.6 11.7 67.6 49.2 0.619 0.685 0.114 0.340 0.336 0.556
Mathpix 0.105 0.384 0.306 0.454 62.7 62.1 77.0 67.1 0.243 0.320 0.108 0.304 0.191 0.365
Docling [30] 0.416 0.987 0.999 1.000 61.3 25.0 0.627 0.810 0.313 0.837 0.589 0.909
Pix2Text [12] 0.138 0.356 0.276 0.611 78.4 39.6 73.6 66.2 0.584 0.645 0.281 0.499 0.320 0.528
Unstructured-0.17.2 0.198 0.481 0.999 1.000 - - - - 1.000 0.998 0.145 0.387 0.586 0.716
OpenParse-0.7.0 0.681 0.974 0.996 1.000 - - 64.8 27.5 0.284 0.639 0.595 0.641 0.646 0.814
Based on Expert VLMs
GOT-OCR [59] 0.189 0.315 0.360 0.528 74.3 45.3 53.2 47.2 0.459 0.520 0.141 0.280 0.287 0.411
Nougat [6] 0.365 0.998 0.488 0.941 15.1 16.8 39.9 0.0 0.572 1.000 0.382 0.954 0.452 0.973
Mistral OCR 0.072 0.325 0.318 0.495 64.6 45.9 75.8 63.6 0.600 0.650 0.083 0.284 0.268 0.439
OLMOCR-sglang 0.097 0.293 0.455 0.655 74.3 43.2 68.1 61.3 0.608 0.652 0.145 0.277 0.326 0.469
SmolDocling-256M 0.262 0.838 0.753 0.997 32.1 55.1 44.9 16.5 0.729 0.907 0.227 0.522 0.493 0.816
Based on General VLMs
GPT-4o [1] 0.144 0.409 0.425 0.606 72.8 42.8 72.0 62.9 0.234 0.329 0.128 0.251 0.233 0.399
Qwen2-VL-72B [55] 0.096 0.218 0.404 0.487 82.2 61.2 76.8 76.4 0.387 0.408 0.119 0.193 0.252 0.327
InternVL2-76B [8] 0.353 0.290 0.543 0.701 67.4 44.1 63.0 60.2 0.547 0.555 0.317 0.228 0.440 0.443
Qwen2.5-VL-7B [4] 0.222 0.236 0.394 0.587 68.4 45.1 71.5 75.7 0.465 0.294 0.246 0.207 0.332 0.331
InternVL3-8B [75] 0.315 0.345 0.714 0.729 59.4 72.3 59.0 71.5 0.352 0.211 0.324 0.257 0.426 0.385
Based on Reinforcement Learning
[…]
3.1 Evaluation and Metrics
OmniDocBench [36]
We introduces an end-to-end evaluation pipeline to ensure fair and unified assessment of document parsing models. It processes model-generated markdown by cleaning, extracting key elements (tables, formulas, code), standardizing inline formulas, and recording positions for reading order evaluation. This comprehensive pipeline enables reliable, multi-level evaluation across diverse document types. We evaluate different content types using tailored metrics :
  • Pure Text: Normalized edit distance (NED) is computed and averaged at the sample level to obtain the final scores.
  • Tables: Tables are converted to HTML, then evaluated using TEDS [71] and NED.
  • Formulas: Evaluation is based on CDM [51], NED, and BLEU.
  • Reading Order: Evaluated by NED, considering only text and excluding other components.
Fox [26]
A multilingual benchmark for fine-grained, text-centric document understanding, covering 9 sub-tasks such as OCR, translation, summarization, layout analysis, and captioning. It includes 212 bilingual pages (112 English, 100 Chinese) with diverse layouts, evaluated using task-specific metrics like NED, F1, BLEU, METEOR, and ROUGE-L.
PubTabNet [71]
A widely used benchmark for table recognition, containing 500,777 training and 9,115 validation images with diverse scientific table structures. Evaluation is conducted on the validation set.
FinTabNet [70]
Focused on financial documents, this dataset includes 112,000 single-page scanned documents, with 92,000 cropped training images and 10,656 for testing. It features dense layouts and detailed annotations for both structure and content evaluation.
olmOCR-Bench [39]
Model Language Table Frame Type Special Situation (+/-) Overall $\uparrow$
EN ZH Mixed Full Omission Three Zero MergeCell Form. Colorful Rotate
PaddleOCR [22] 76.8 71.8 80.1 67.9 74.3 81.1 74.5 70.6/75.2 71.3/74.1 72.7/74.0 23.3/74.6 73.6
RapidTable [40] 80.0 83.2 91.2 83.0 79.7 83.4 78.4 77.1/85.4 76.7/83.9 77.6/84.9 25.2/83.7 82.5
StructEqTable [74] 72.8 75.9 83.4 72.9 76.2 76.9 88.0 64.5/81.0 69.2/76.6 72.8/76.4 30.5/76.2 75.8
GOT-OCR [59] 72.2 75.5 85.4 73.1 72.7 78.2 75.7 65.0/80.2 64.3/77.3 70.8/76.9 8.5/76.3 74.9
Qwen2-VL-7B [55] 70.2 70.7 82.4 70.2 62.8 74.5 80.3 60.8/76.5 63.8/72.6 71.4/70.8 20.0/72.1 71.0
InternVL2-8B [8] 70.9 71.5 77.4 69.5 69.2 74.8 75.8 58.7/78.4 62.4/73.6 68.2/73.1 20.4/72.6 71.5
Qwen2.5-VL-7B [55] 87.4 80.7 93.5 86.4 85.1 84.1 88.7 77.5/89.8 82.1/87.2 77.1/87.5 56.5/86.0 85.5
InternVL3-8B [75] 79.5 86.0 91.7 85.5 80.7 83.9 85.9 71.9/90.9 74.0/86.7 82.1/85.3 12.6/85.5 84.3
[…]
3.2 Main Results
In this section, we evaluate the model’s performance in document parsing, table recognition, and Document-level OCR. In this work, we ensure that the test data for each benchmark undergoes rigorous text similarity filtering to prevent any overlap with the training data.
Overall Evaluation on OmniDocBench As shown in Table 2, pipeline-based methods such as MinerU [52] and Mathpix achieve superior performance across individual sub-tasks including text recognition, formula recognition, and table recognition. Meanwhile, general-purpose vision-language models like Qwen2-VL and GPT-4o also demonstrate competitive results. Notably, most methods perform better on English pages compared to Chinese pages, reflecting language-dependent challenges. In contrast, our proposed Infinity-Parser-7B achieves a more balanced performance across all sub-tasks and languages, setting new SOTA results in table recognition and maintaining competitive accuracy in text, formula, and reading order extraction. This highlights the advantage of reinforcement learning with multi-aspect rewards in enabling robust, end-to-end document parsing.
Table Recognition Evaluation To evaluate the model’s generalization ability, we introduce task-specific test cases. In Table 4, we compare Infinity-Parser-7B with end-to-end table recognition models on PubTabNet and FinTabNet using the TEDS metric, which evaluates both structure and content. We also report TEDS-S for structure-only assessment. The evaluation results for InternVL3, Qwen2.5-VL, and GPT-4o were generated through our standardized benchmarking pipeline. Infinity-Parser-7B achieves the highest TEDS-S and TEDS scores on both datasets.
Model PubTabNet [71] FinTabNet [70]
TEDS-S TEDS TEDS-S TEDS
EDD [72] 89.9 88.3 90.6 -
OmniParser [49] 90.45 88.83 91.55 89.75
InternVL3-8B [75] 87.48 83.02 86.73 84.01
InternVL3-78B [75] 89.63 82.11 92.51 89.21
Qwen2.5-VL-7B [3] 86.78 81.60 87.46 82.58
Qwen2.5-VL-72B [3] 87.91 84.39 87.13 82.90
GPT-4o [1] 86.16 76.53 87.00 83.96
[…]
Document-level OCR Evaluation As shown in Table 5, Infinity-Parser-7B achieves the best edit distance among all evaluated models, scoring 0.023 for English and 0.017 for Chinese, outperforming all baselines including large-scale models such as Qwen-VL-Max [2] and Fox [26]. This result suggests that Infinity-Parser-7B can reconstruct dense document text with exceptional precision and minimal character-level errors.
@@ 3.2 Main Results
UReader [67] 7B 0.344 - 0.296 - 0.469 - 0.103 - 0.287 - 0.718 -
LLaVA-NeXT [27] 34B 0.647 - 0.573 - 0.881 - 0.478 - 0.582 - 0.430 -
InternVL-ChatV1.5 [7] 26B 0.751 0.816 0.698 0.784 0.917 0.866 0.568 0.622 0.663 0.717 0.393 0.265
Nougat [5] 250M 0.745 - 0.720 - 0.809 - 0.665 - 0.761 - 0.255 -
TextMonkey [28] 7B 0.821 - 0.778 - 0.906 - 0.671 - 0.762 - 0.265 -
DocOwl1.5 [14] 7B 0.862 - 0.835 - 0.962 - 0.788 - 0.858 - 0.258 -
Vary [57] 7B 0.918 0.952 0.906 0.961 0.956 0.944 0.885 0.754 0.926 0.873 0.092 0.113
Vary-toy [58] 1.8B 0.924 0.914 0.919 0.928 0.938 0.907 0.889 0.718 0.929 0.832 0.082 0.142
Qwen-VL-Plus [2] - 0.931 0.895 0.921 0.903 0.950 0.890 0.893 0.684 0.936 0.828 0.096 0.121
[…]
Fox [26] 1.8B 0.952 0.954 0.957 0.964 0.948 0.946 0.930 0.842 0.954 0.908 0.046 0.061
GOT [60] 580M 0.972 0.980 0.971 0.982 0.973 0.978 0.947 0.878 0.958 0.939 0.035 0.038
[…]
Table 5: Performance comparison of dense English and Chinese OCR on document-level pages.
Models Book Slides Financial Report Textbook Exam Paper Magazine Academic Papers Notes Newspaper Overall $\downarrow$

### v3

@@ 3.2 RL with Layout-Aware Rewards
As illustrated in Figure 3, we employ an RL framework to directly optimize scanned document parsers, aiming to enhance both structural fidelity and semantic accuracy. Specifically, we utilize GRPO (Shao et al., 2024), which enables learning from rule-based reward signals without relying on absolute values. GRPO operates by generating a set of candidate Markdown outputs for each document and evaluating them using a multi-aspect reward, denoted as $R_{\text{Multi-Aspect}}$, which integrates multiple rule-based criteria into a unified supervisory signal. These raw rewards are then converted into relative advantage scores by comparing each candidate against others within the same group. This relative evaluation promotes training stability and encourages the selection of higher-quality outputs, eliminating the need for a learned value function or critic. Notably, our reinforcement learning approach avoids any explicit thinking or intermediate reasoning process; instead, all outputs are treated as final answers, with the model receiving verifiable rewards based on these outputs.
@@
Edit Distance Reward ($R_{\mathrm{dist}}$)
We define the edit distance reward based on the normalized Levenshtein distance D(y,y^)D($\hat{y}$) between the predicted output $\hat{y}$ and the reference output $y$:
$R_{\text{dist}}=1-\tfrac{D(y,\hat{y})}{\max(N,M)}$ (1)
[…]
Figure 3: Overview of Infinity-Parser training framework. Our model is optimized via reinforcement finetuning with edit distance, layout, and order-based rewards.
Count Reward ($R_{\mathrm{count}}$)
@@
$R_{\text{count}}=1-\tfrac{|N_{Y}-N_{\hat{Y}}|}{N_{Y}}$ (2)
@@
Order Reward ($R_{\mathrm{order}}$)
@@
$R_{\text{order}}=1-\frac{D_{\text{order}}}{\max_{\text{inv}}}$ (3)
@@
The final multi-aspect reward is a weighted combination of these three components. Specifically, we begin by applying the Hungarian algorithm (Kuhn, 1955) to establish the optimal one-to-one matching between predicted and ground-truth segments, identifying both pairings and their relative order. Based on the matched segment count, we compute the count reward to reflect alignment in the number of segments. Using the relative sequence of matched pairs, we calculate the order reward to measure structural consistency. On top of these matchings, we compute the edit reward by averaging the edit similarities of each matched segment pair. Combining these terms yields the final reward:
$R_{\text{Multi-Aspect}}=R_{\text{dist}}+R_{\text{count}}+R_{\text{order}}$ (4)
@@ 4.1 Implementation Details
We fine-tune the Qwen2.5-VL-7B model using GRPO within a distributed training setup based on Verl Sheng et al. (2024); Yaowei Zheng (2025), utilizing 8 A100 GPUs (80GB). Throughout our experiments, we set the KL coefficient $\beta=1.0\times 10^{-2}$. And for each problem instance, we sample 8 responses, each with a maximum length of 8192 tokens and a temperature of 1.0. Both the rollout batch size and the global batch size are set to 128. The actor model is updated using the AdamW optimizer with parameters $(\beta_{1}=0.9,\beta_{2}=0.99)$ and a learning rate $1.0\times 10^{-6}$. The model is trained for 1.0 epoch for all experiments. Due to limited computational resources, we randomly sampled 43K documents from the 400K corpus for training. In our main results, we directly performed reinforcement learning on the base model using the 43K subset.
@@
Based on Pipeline Tools
MinerU (Wang et al., 2024b) 0.15 0.357 0.061 0.215 0.278 0.577 78.6 62.1 0.18 0.344 0.079 0.292
Marker (Paruchuri, 2024) 0.336 0.556 0.080 0.315 0.530 0.883 67.6 49.2 0.619 0.685 0.114 0.340
Mathpix 0.191 0.365 0.105 0.384 0.306 0.454 77.0 67.1 0.243 0.320 0.108 0.304
Docling (Livathinos et al., 2025) 0.589 0.909 0.416 0.987 0.999 1.000 61.3 25.0 0.627 0.810 0.313 0.837
Pix2Text (Gurgurov & Morshnev, 2024) 0.320 0.528 0.138 0.356 0.276 0.611 73.6 66.2 0.584 0.645 0.281 0.499
Unstructured-0.17.2 0.586 0.716 0.198 0.481 0.999 1.000 - - 1.000 0.998 0.145 0.387
OpenParse-0.7.0 0.646 0.814 0.681 0.974 0.996 1.000 64.8 27.5 0.284 0.639 0.595 0.641
Based on Expert VLMs
GOT-OCR (Wei et al., 2024) 0.287 0.411 0.189 0.315 0.360 0.528 53.2 47.2 0.459 0.520 0.141 0.280
Nougat (Blecher et al., 2024) 0.452 0.973 0.365 0.998 0.488 0.941 39.9 0.0 0.572 1.000 0.382 0.954
Mistral OCR 0.268 0.439 0.072 0.325 0.318 0.495 75.8 63.6 0.600 0.650 0.083 0.284
OLMOCR-sglang 0.326 0.469 0.097 0.293 0.455 0.655 68.1 61.3 0.608 0.652 0.145 0.277
SmolDocling-256M 0.493 0.816 0.262 0.838 0.753 0.997 44.9 16.5 0.729 0.907 0.227 0.522
Based on General VLMs
GPT-4o (Achiam et al., 2023) 0.233 0.399 0.144 0.409 0.425 0.606 72.0 62.9 0.234 0.329 0.128 0.251
Qwen2-VL-72B (Wang et al., 2024c) 0.252 0.327 0.096 0.218 0.404 0.487 76.8 76.4 0.387 0.408 0.119 0.193
InternVL2-76B (Chen et al., 2024) 0.440 0.443 0.353 0.290 0.543 0.701 63.0 60.2 0.547 0.555 0.317 0.228
Qwen2.5-VL-7B (Bai et al., 2025) 0.220 0.265 0.142 0.205 0.393 0.530 78.7 78.3 0.155 0.162 0.191 0.169
InternVL3-8B (Zhu et al., 2025) 0.426 0.385 0.315 0.345 0.714 0.729 59.0 71.5 0.352 0.211 0.324 0.257
Based on Reinforcement Learning
Infinity-Parser-7B 0.141 0.197 0.076 0.117 0.314 0.434 85.3 81.4 0.098 0.142 0.076 0.095
4.2 Main Results
We evaluate our method on several widely-used benchmarks for document understanding and OCR tasks. OmniDocBench (Ouyang et al., 2024) provides comprehensive evaluation across diverse document types using NED and TEDS metrics. For table recognition, we use PubTabNet with scientific tables and FinTabNet (Zheng et al., 2021) with financial documents. Additionally, we employ olmOCR-Bench (Poznanski et al., 2025) for fact-based OCR evaluation. We ensure that the test data for each benchmark undergoes rigorous text similarity filtering to prevent any overlap with the training data. Detailed descriptions of these benchmarks are provided in Appendix.
Overall Evaluation on OmniDocBench
As shown in Table 2, pipeline-based methods such as MinerU (Wang et al., 2024b) and Mathpix achieve superior performance across individual sub-tasks including text recognition and formula recognition. Meanwhile, general-purpose vision-language models like Qwen2.5-VL-7B and GPT-4o also demonstrate competitive results. Notably, most methods perform better on English pages compared to Chinese pages, reflecting language-dependent challenges. In contrast, our proposed Infinity-Parser-7B achieves a more balanced performance across all sub-tasks and languages, setting new SOTA results with overall edit distances of 0.141 and 0.197. This highlights the advantage of reinforcement learning with multi-aspect rewards in enabling robust, end-to-end document parsing.
Model Overall ArXiv Old Scans Math Tables Old Scans Headers&Footers Multi Col. Long-Tiny Text Base
GOT OCR 48.3 52.7 52.0 0.2 22.1 93.6 42.0 29.9 94.0
Marker v1.6.2 59.4 24.3 22.1 69.8 24.3 87.1 71.0 76.9 99.5
MinerU v1.3.10 61.5 75.4 47.4 60.9 17.3 96.6 59.0 39.1 96.6
Mistral OCR API 72.0 77.2 67.5 60.6 29.3 93.6 71.3 77.1 99.4
GPT-4o (No Anchor) 68.9 51.5 75.5 69.1 40.9 94.2 68.9 54.1 96.7
GPT-4o (Anchored) 69.9 53.5 74.5 70.0 40.7 93.8 69.3 60.6 96.8
Gemini Flash 2 (No Anchor) 57.8 32.1 56.3 61.4 27.8 48.0 58.7 84.4 94.0
Gemini Flash 2 (Anchored) 63.8 54.5 56.1 72.1 34.2 64.7 61.5 71.5 95.6
Qwen 2 VL (No Anchor) 31.5 19.7 31.7 24.2 17.1 88.9 8.3 6.8 55.5
Qwen 2.5 VL (No Anchor) 65.5 63.1 65.7 67.3 38.6 73.6 68.3 49.1 98.3
olmOCR v0.1.68 (No Anchor) 76.3 72.1 74.7 71.5 43.7 91.6 78.5 80.5 98.1
olmOCR v0.1.68 (Anchored) 77.4 75.6 75.1 70.2 44.5 93.4 79.4 81.7 99.0
Infinity-Parser-7B 82.5 84.4 83.8 85.0 47.9 88.7 84.2 86.4 99.8
Table 3: Performance comparison on the olmOCR (Poznanski et al., 2025) benchmark across multiple document domains and structural challenges. Higher is better.
Document-level OCR Evaluation Table 3 reports performance on the olmOCR-Bench benchmark, which evaluates document-level OCR across diverse layouts and domains. Infinity-Parser-7B achieves the highest overall score (82.5), followed closely by olmOCR v0.1.68 (Anchored) (77.4), both demonstrating strong performance in complex categories like multi-column layouts and scanned math content. The results highlight the effectiveness of anchored prompting, with anchored versions of models (e.g., GPT-4o, olmOCR) significantly outperforming their non-anchored counterparts—especially on tables and old scans. This underscores the importance of layout-aware extraction techniques. In contrast, traditional pipelines like Marker and GOT OCR lag behind in structural accuracy, reinforcing the value of modern VLM-based approaches in high-fidelity PDF understanding.
Table Recognition Evaluation
To evaluate the model’s generalization ability, we introduce task-specific test cases. In Table 4, we compare Infinity-Parser-7B with end-to-end table recognition models on PubTabNet and FinTabNet using the TEDS metric, which evaluates both structure and content. We also report TEDS-S for structure-only assessment. The evaluation results for InternVL3, Qwen2.5-VL, and GPT-4o were generated through our standardized benchmarking pipeline. Infinity-Parser-7B achieves the highest TEDS-S and TEDS scores on both datasets.
Model PubTabNet FinTabNet
TEDS-S TEDS TEDS-S TEDS
EDD 89.9 88.3 90.6 -
OmniParser 90.45 88.83 91.55 89.75
InternVL3-8B 87.48 83.02 86.73 84.01
InternVL3-78B 89.63 82.11 92.51 89.21
Qwen2.5-VL-7B 86.78 81.60 87.46 82.58
Qwen2.5-VL-72B 87.91 84.39 87.13 82.90
GPT-4o 86.16 76.53 87.00 83.96
Infinity-Parser-7B 93.46 91.82 97.16 95.92
4.3 Ablation Study
We perform ablation experiments to evaluate the individual contributions of our three core design choices: (1) data quality verification and (2) multi-aspect rewards. We report all ablation results using two primary evaluation metrics: OverallEdit and OverallCat.. OverallEdit represents the average edit-based overall score across English and Chinese pages, as shown in Table 2. In contrast, OverallCat. reflects the mean category-level performance across nine types of scanned document pages, following the same evaluation setting as Table 9 in the Appendix.
Method Edit Dist. Count. Order. SFT RL Overall (EN)Edit $\downarrow$ Overall (ZH)Edit $\downarrow$ Overall Cat. $\downarrow$
Zero Shot - - - - - 0.220 0.265 0.183
SFT - - - 43K - 0.198 0.261 0.159
Zero + RL - - - 43K 0.169 0.224 0.156
Zero + RL - - 43K 0.159 0.200 0.112
Zero + RL - 43K 0.141 0.197 0.104
SFT + RL 43K 43K 0.163 0.195 0.092
Table 5: Results under different reward designs.
Effect of Multi-Aspect Rewards.
Table 5 demonstrates that reinforcement learning can outperform supervised fine-tuning when appropriate reward designs are applied. Compared to the SFT baseline (0.198 / 0.261 / 0.159), the RL method with distance-based reward (Zero + $R_{\text{dist}}$) achieves better OverallEdit (EN: 0.169 vs. 0.198, ZH: 0.224 vs. 0.261) while maintaining a comparable OverallCat. (0.156 vs. 0.159). Incorporating additional count and order rewards further improves structural consistency: Zero + $R_{\text{dist}}+R_{\text{count}}$ achieves 0.159 / 0.200 / 0.112, and Zero + $R_{\text{dist}}+R_{\text{count}}+R_{\text{order}}$ achieves 0.141 / 0.197 / 0.104. Meanwhile, when reinforcement learning is combined with supervised fine-tuning (SFT + RL), the model does not exhibit further significant improvements. This observation is consistent with existing studies (Guo et al., 2025), which suggest that given the backbone model’s inherent capabilities, SFT may be unnecessary for downstream RL training in certain scenarios. These results further demonstrate that reinforcement learning, when equipped with structural supervisory signals, enables the model to better align with task-specific objectives.
Figure 4: Performance comparison of SFT and Layout-Aware RL on OmniDocBench sub-tasks.
Figure 5: Comparison of model performance on different document parsering tasks.
4.4 Further Analysis of LayoutRL
We analyze the behavior of Layout-Aware RL compared with SFT across various document understanding settings. Our analysis focuses on three aspects: training stability, robustness across diverse document types, and generalization under distribution shifts. In particular, the third analysis examines generalization when the training data do not include the textbook and slides domains, which are used only for evaluation. This setup naturally forms OOD scenarios, allowing us to assess how well each model generalizes to unseen document types and layouts.
Training Stability Across Task Types.
As shown in Figure 4, we compare SFT and Layout-Aware RL on four OmniDocBench sub-tasks: table recognition, text recognition, formula parsing, and reading order prediction. Across all tasks, RL consistently achieves better performance, yielding higher scores on TEDS and lower errors on NED. More importantly, the RL curves exhibit smoother trajectories with stable improvements over training, while SFT shows large fluctuations and even performance regressions at certain stages. These results indicate that layout-aware rewards not only improve final accuracy but also enhance training stability throughout optimization.
Figure 6: Generalization comparison of SFT and Layout-Aware RL. X-axis represents training steps. Y-axis represents 1-NED scores.
Robustness Across Diverse Document Tasks.
Figure 5 compares RL, SFT, and Zero-Shot (Base) across diverse document parsing tasks. On olmOCR (left), RL consistently achieves higher Levenshtein distance similarity score, especially on challenging cases like old scans and table tasks. At the same time, SFT offers moderate gains over Zero-Shot but remains behind RL. On OmniDocBench (right), RL also outperforms the other methods across most document types, showing notable improvements on books, reports, and academic texts. Overall, RL demonstrates greater robustness and better generalization in both structural parsing on olmOCR and text recognition on omnidocbench.
Analysis of Generalization
The Figure 6 compares document parsing performance across different training steps. The top two plots (magazine, research report) correspond to In-Distribution settings, where training and evaluation domains are aligned, while the bottom two plots (colorful textbook, slides) correspond to OOD settings, where evaluation involves unseen document types. In the in-distribution case, SFT achieves stable paragraph-level accuracy but its page-level performance tends to plateau, reflecting reliance on surface patterns. By contrast, RL continues to improve with data scale, achieving notable gains in page-level accuracy. In the OOD case, SFT performance degrades more severely, while RL maintains robustness and shows stronger improvements, highlighting its ability to capture global structural dependencies and generalize across distributions.
5 Conclusion
We introduced LayoutRL, an end-to-end reinforcement learning framework that explicitly incorporates layout awareness into document parsing through verifiable, multi-aspect rewards. To support this training, we built Infinity-Doc-400K, a large-scale dataset combining synthetic and real-world documents with diverse layouts, and trained Infinity-Parser, a VLM-based parser. Experiments on OmniDocBench, olmOCR-Bench, PubTabNet, and FinTabNet show that our approach achieves state-of-the-art performance across languages and document types, outperforming both specialized pipelines and general-purpose VLMs. Beyond accuracy, LayoutRL improves training stability and demonstrates robustness across diverse document tasks, highlighting reinforcement learning as a promising direction for robust and transferable document intelligence.
Appendix A Reproducibility Statement
We are committed to ensuring the reproducibility of our results. To this end, we will release the full Infinity-Doc-400K dataset, the implementation of LayoutRL with verifiable reward functions, and the pretrained Infinity-Parser model. Detailed training configurations, hyperparameters, and evaluation protocols are provided in the appendix and supplementary materials. Our experiments are conducted on widely used benchmarks, including OmniDocBench, olmOCR-Bench, PubTabNet, and FinTabNet, ensuring comparability with prior work. We will also provide scripts for preprocessing, training, and evaluation to facilitate reproducibility and further research by the community.
Appendix B Training Details
B.1 Training Context Length Distribution
Our model was trained with a maximum context length of 8K tokens. To provide a clearer picture of the training data, we report detailed statistics of the context length distribution in Table 6 and Table 7. The average context length is 1,765 tokens, with a maximum of 31,147 tokens. More than 73% of the samples fall within the [512, 4K) range. For sequences exceeding the 8K limit, we applied a left-truncation strategy to retain the semantically more relevant content at the end of the sequence.
Metric Min Max Average Median Std
Value 17 31,147 1,765 1,127 1,692
Table 6: Summary statistics of training context length.
Context Length (tokens) [1,256) [256,512) [512,1K) [1K,2K) [2K,4K) [4K,8K) [8K,16K) $\geq$16K
Frequency (count) 25,125 40,680 119,041 102,757 70,554 39,780 2,482 63
Distribution (%) 6.27 10.16 29.72 25.66 17.62 9.93 0.62 0.02
Table 7: Distribution of training samples across different context length intervals.
Appendix C Data Details
As illustrated in Figure 7, the dataset spans seven diverse document domains, making it one of the most richly annotated and structurally varied resources to date. Each domain is represented by two sample pages, highlighting the broad variability in layout design, content structure, and semantic density. For instance, Medical Reports typically contain structured tables with clinical measurements and diagnostic notes. Synthetic Documents are algorithmically generated to replicate real-world formats, providing layout diversity for training robust parsers. Financial Reports feature dense tables and formal accounting records, while Academic Papers often follow two-column layouts with references, equations, and figures. Books combine narrative content with visual illustrations, and Magazines blend images and stylized text for reader engagement. Finally, Web Pages, when saved as PDFs, preserve HTML-based structures that integrate tables, lists, and dynamic elements. This visual taxonomy exemplifies the structural and semantic diversity present in real-world documents, highlighting the core challenge faced by document AI systems: reliably parsing heterogeneous layouts and extracting structured information across a wide variety of formats.
Table 8 provides an overview of the document types included in the Infinity-Doc-400K dataset, detailing the composition across both real-world and synthetic sources. The real-world portion consists of 331K documents spanning six domains: financial reports, medical reports, academic papers, books, magazines, and web pages. These documents were collected from the web and annotated using a pseudo-labeling pipeline based on expert model agreement. While this approach enables large-scale data acquisition, the resulting label quality is relatively low due to occasional inconsistencies across models. Additionally, real-world data collection incurs a high cost, especially in terms of manual filtering, formatting normalization, and layout validation.
Data Source Document Types Size Annotation Method Cost
Real-World Doc Financial Reports 58.0K Web + Pseudo-Label High
Real-World Doc Medical Reports 5.0K Web + Pseudo-Label High
Real-World Doc Academic Papers 71.7K Web + Pseudo-Label High
Real-World Doc Books 11.3K Web + Pseudo-Label High
Real-World Doc Magazines 180.0K Web + Pseudo-Label High
Real-World Doc Web Pages 5.0K Web + Pseudo-Label High
Synthetic Synthetic Documents 69.0K CC3M + Web + Wiki Low
Table 8: Overview of document types in the Infinity-Doc-400K dataset, including data source, document type, annotation method, and collection cost.
Appendix D Benchmarks Details
OmniDocBench (Ouyang et al., 2024) We conduct evaluation on OmniDocBench, a comprehensive benchmark that covers diverse document types and content modalities. To assess parsing performance across different structural elements, we employ two primary evaluation metrics: Normalized Edit Distance (NED), which measures the minimum edit operations required to transform one string into another normalized by the target string length, and Tree Edit Distance-based Similarity (TEDS), which captures structural similarities by comparing tree representations of HTML tables. These metrics are applied to different subtasks: NED is used to evaluate pure text, formula transcription, and reading order; TEDS combined with NED is used to evaluate both structural and content accuracy of table parsing.
PubTabNet (Zhong et al., 2020) A widely used benchmark for table recognition, containing 500,777 training and 9,115 validation images with diverse scientific table structures. Evaluation is conducted on the validation set.
FinTabNet (Zheng et al., 2021) Focused on financial documents, this dataset includes 112,000 single-page scanned documents, with 92,000 cropped training images and 10,656 for testing. It features dense layouts and detailed annotations for both structure and content evaluation.
olmOCR-Bench (Poznanski et al., 2025)
Appendix E More Results
Diverse Page Types Evaluation
To further investigate model behavior across diverse document types, we evaluated text recognition performance on nine distinct page categories. As shown in Table 9, pipeline-based systems such as MinerU Wang et al. (2024b) and Mathpix achieved strong results on structured formats like academic papers and financial reports. General-purpose vision-language models (VLMs) demonstrated better generalization on less formal page types, including presentation slides and handwritten notes. However, for challenging formats such as newspapers, most VLMs underperformed, while pipeline tools maintained relatively lower error rates. Notably, our proposed Infinity-Parser-7B achieved consistently low edit distances across all document types, outperforming both pipeline-based systems and general-purpose VLMs in overall accuracy. This highlights the robustness and adaptability of our reinforcement learning approach across diverse and complex document layouts.
Models Book Slides
Financial
Report
Textbook
Exam
Paper
Magazine
Academic
Papers
Notes Newspaper Overall $\downarrow$